A double slit interference pattern is produced by a parallel beam of light passing
through two slits of unequal widths 20 λ and 40 λ, respectively, where λ is the
wavelength of the light used. The distance between the centres of the slits is
1,000 λ. If the fringes are observed on a screen far off from the slits (D >> 1,000 λ),
calculate the separation of two adjacent maxima.
1
Expert's answer
2019-03-25T08:52:53-0400

https://en.m.wikipedia.org/wiki/Double-slit_experiment

"w = D\frac{\lambda}{d} = D\frac{\lambda}{1000\lambda} = \frac{D}{1000}."